OpenStudy (hang254):

Find the first six terms of the sequence. a1 = -4, an = an-1 + 7

OpenStudy (solomonzelman):

\(\large\color{blue}{ a_1=-4 }\) saying that the 1st term is `-4`. \(\large\color{blue}{ a_n=a_{n-1}+7 }\), this means that the difference is equal to `7`.

OpenStudy (hang254):

ok, so would i add 7 after -4 and so on?

OpenStudy (solomonzelman):

\(\large\color{blue}{ a_n=a_1 + d(n-1) }\), and in your case, \(\large\color{blue}{ a_5=a_1 + d(5-1) }\) since looking for the fifth term. Plug in the `a1` and the difference that you know., and you get \(\large\color{blue}{ a_5=-4 + 7(5-1) }\)

OpenStudy (solomonzelman):

So, \(\large\color{black}{ a_5=??? ~~~~~~~\rm{(you~~~tell~~~me)} }\)

OpenStudy (hang254):

12

OpenStudy (solomonzelman):

\(\large\color{blue}{ a_5=-4 + 7(5-1) }\) \(\large\color{blue}{ a_5=-4 + 7(4) }\) \(\large\color{blue}{ a_5=-4 +28 }\) \(\large\color{blue}{ a_5=? }\) (not 12)

OpenStudy (hang254):

omg i forgot to do PEMDAS, lol sorry 24

OpenStudy (hang254):

that makes more sense now

OpenStudy (hang254):

so it would be: -4, 3, 10, 17, 24, 31 for the first 6 terms